OpenStudy (anonymous):

compute the sum for other telescoping series : ln 1/2 + ln 2/3 + 3/4

OpenStudy (experimentx):

or - infinity

OpenStudy (callisto):

I assume you're saying ln 1/2 + ln 2/3 + ln3/4 ln 1/2 + ln 2/3 + ln3/4 = (ln 1- ln2)+(ln2-ln3) + (ln3-ln4) = ln1 - ln4 = 0-ln4 =-ln4 If it is like ln 1/2 + ln 2/3 + ln3/4 + .... + ln (n-1)/n then it is ln1 - ln (n) = -ln (n) If it is sum to infinity then it is -infinity

OpenStudy (anonymous):

help please Boxes of car parts are stacked on top of each other on a work bench. The table below shows how the height above the floor of the topmost box depends on the number of boxes. What is a rule for the height? Give the rule in words and as an algebraic expression. Number of Boxes Height (in.) 2 (9 • 2) + 39 3 (9 • 3) + 39 4 (9 • 4) + 39 n ?

OpenStudy (anonymous):

@elena if you want someone to answer your question it is easier to have it posted on the forum rather than my question

OpenStudy (anonymous):

i did and no body helped

OpenStudy (anonymous):

Bump ur question, @elena

OpenStudy (anonymous):

i did so many times